### v2

@@ 2.3 Oscillations in Auto-Differentiation
To start, we examine Leapfrog. Comparing (11) and (15), we immediately observe some nice features. $\partial_{\theta_{l}}E$ honors the structure of $\frac{\delta\mathcal{E}}{\delta\theta}(lh)$. Both of them see the gradient as a product of $p$, the cotangent vector that runs from final time to $t=0$, and the Jacobian $\partial_{\theta}f$. Moreover, the cotangent vector of $\mathsf{p}$ aligns very well with that of (12). In particular, (16c) is exactly a Leapfrog discretization of (12) for all $l\leq L-3$.
However, one strong drawback of the formula hinges on the definition of $\mathsf{p}_{L-1}$ an $\mathsf{p}_{L-2}$. Intuitively, for a good match, both $\mathsf{p}_{L-1}$ and $\mathsf{p}_{L-2}$ should approximately resemble $p(1)$. Comparing (16a) and (12), however, we see that $\mathsf{p}_{L-1}$ nearly doubles the value of $p(1)$ while $\mathsf{p}_{L-2}$ is $\mathcal{O}(h)$. This immediately introduces $\mathcal{O}(1)$ contrast at the beginning of the evolution of $\mathsf{p}$. This high contrast propagates back in time using the Leapfrog formula (16c), bringing high oscillations throughout. This was seen in Figure 1. Moreover, refining the discretization by enlarging $L$ and reducing $h$ is not helpful. The $\mathcal{O}(1)$ oscillation is encoded in the definition of $\mathsf{p}_{L-1}$ and $\mathsf{p}_{L-2}$, and it does not disappear in the $h\to 0$ limit.
We now examine ODE-Net generated by ERK-2 stage method. The situation here is even more intricate. This time, one can show that the cotangent vector $\mathsf{p}_{l}$ (18) is actually a second order approximation of $p$:
@@
The proof for this proposition is straightforward, as presented in Supplementary Section S3.2. Essentially, one needs to define an auxiliary variable, termed $\hat{\mathsf{p}}_{l}$, that solves the same equation (18) with $\mathsf{z}_{l}$ and $\mathsf{z}_{l+\alpha}$ replaced by $z(lh)$ and $z((l+\alpha)h)$ respectively, and deploys the triangular inequality.
@@
Higher order ERK suffers from the same challenge. Though the assembling coefficient vary, the typical situation nevertheless gives incorrect match to its continuous versions, and provides high oscillations. For example, in the case of Nystrom (8),
@@ 3 Modified auto-differentiation for ODE-net
The methods provided below require some generic assumptions for the fix to hold. {assumption} The activation function $f(z,\theta)$ satisfies: a) Regularity: $f$ is third order continuously differentiable with derivatives bounded by $C_{a}$; b) Quadratic-linear growth: For all $z\in\mathbb{R}^{d},\theta\in\mathbb{R}^{n}$, $|f|\leq C_{b}(|\theta|^{2}+1)(|z|+1)$ for some $C_{b}$. The output function $g(z)$ to possess: c) Regularity: $g(z)$ and its gradient $\nabla g(z)$ are Lipschitz continuous with Lipschitz constants bounded by $C_{c}$. The training data distribution $\mu$ has: d) Compact Support: There exists some $C_{d}>0$ such that $\text{supp}(\mu)$, the support of $\mu$, lies within a ball centered at the origin with radius $C_{d}$. These assumptions are rather mild and are typically satisfied. For example, the commonly used activation function $f(z;\sigma,W,b)=\sigma\cdot\tanh(Wz+b)$ and the output function $g(z):=z$ satisfiy all the conditions above.
@@ 3.1 Modified Leapfrog ODE-net
The core of the proof lies in the realization that “averaging” the gradient is effectively equivalent (up to $\mathcal{O}(h^{2})$ error) to “averaging” the cotangent vector $\mathsf{p}$. The averaging for the gradient is already denoted by $\mathsf{T}$ defined in (21), and we are to define a new matrix $\tilde{\mathsf{T}}$ (see definition in (S(22))) that acts on $\{\mathsf{p}\}$. Two auxiliary quantities will be defined. Firstly, we define $\hat{\mathsf{p}}=\{\hat{\mathsf{p}}_{l}\}_{l=0}^{L-1}$, a semi-continuous object that uses the same formula (16) with $\mathsf{z}_{l}$ replaced by $z(lh)$. We then define an averaged version:
@@
The theorem is proved if all three terms are $\mathcal{O}(h^{2})$. This can be proved to hold. We discuss the proof strategy in Appendix A.1, and leave details to supplementary materials in Corollary S2.9, Lemma S2.13 and Corollary S2.11 respectively.
@@ 3.2 Modified 2-stage ERK ODE-nets
We should note that $\{\mathsf{p}_{l}\}$ is an output of auto-differentiation, as justified by the following proposition:
@@
For all $l=0,1,\dots,L$, the solution $\mathsf{p}_{l}$ to  (18) satisfies: $\mathsf{p}_{l}^{\top}=\partial_{\mathsf{z}_{l}}E$.
The proof is straightforward calculation and is in Supplementary Section S3.1.
@@
We leave the discussion of the proof strategy to Appendix A.2, with details provided in Supplementary Section S3. The major component of the proof is Taylor approximation both for the forward and the backward equation to the second order.
@@ 4 Numerical experiments
We present numerical experiments to verify our observations, and validate the correction method we proposed. The implementation details and additional figures are deferred to Appendix B and Supplementary Section S4 respectively.
Four discretization schemes are studied, and they are Leapfrog (3), an example of LLM, Midpoint and Ralston (6)-(7), as examples of ERK-2 stage, and Nystrom (8) as an example of ERK-3 stage respectively. For the first three schemes, we provide corrections.
@@ 4.1 Gradient computation
In Figure 2S1, and S2, we plot the performance of auto-differentiation result and the our proposed correction for Leapfrog, Midpoint and Ralston respectively. In each plot, the left panel shows comparison of true gradient (blue), gradient as an output of auto-differentiation (green) and our modified correction (red), following (22) and (26). The right panel demonstrates the convergence rate with respect to $h$. For all three methods, the output of auto-differentiation are highly oscillatory and non-converging: In $L_{\infty}$ norm, the error stays as a constant for all $h$. The modified gradients (red) visually show no difference from the true gradient, and the convergence rate with respect to $h$ are $h^{2.00}$, $h^{1.70}$, and $h^{1.84}$ respectively, confirming the second order accuracy predicted in Theorem 3.1 and 3.4. As a comparison, we also plot out the convergence using the standard forward Euler discretization. The rate stays around $h^{~0.98}$, and thus confirms it as a first order method.
@@ 4.2 Curve-learning
Figure 3: For Leapfrog Method. The top row shows the training loss and test error in epochs. The middle row compares $\Theta$ outputs from vanilla auto-differentiation method and our modified gradients. The bottom plot presents the final actual reconstruction.
Figure 4: For Leapfrog Method. The figure compares reconstructions of a spiral learned using auto-differentiation gradients (red) and our modified gradients (blue). Our modification achieves results that align more closely with the ground truth (green).
@@
We first experimented on learning a linear function. The improvement for Leapfrog, Midpoint and Ralston are respectively shown in Figure 3S3, and S4. The top rows show both auto-differentiation updates (in red) and our modified gradients (in blue) decay in error over epochs. In these examples, our test error clearly shows better performance compared to auto-differentiation. The second rows depict the computed $\Theta$. While auto-differentiation leads to strong oscillatory patterns, our modified gradients provide smoother parameters and better convergence in epoches. The last rows present the final function reconstructions. For a linear function, our modified gradients show consistent improvement toward the ground truth (in green).
We also tested learning a spiral in $\mathbb{R}^{2}$. The results for Leapfrog, Midpoint, and Ralston methods are shown in Figures 4, S5, and S6, respectively. In each plot, the reconstructed trajectory using auto-differentiation is shown in red, while the reconstruction using the modified gradient is shown in blue. Our modification produces results that more closely match the ground truth, shown in green.
@@ 5 Conclusions
Our findings suggest that careful treatment of discretization and optimization order is crucial for improving stability and accuracy in training physics-inspired neural networks. Future work may explore extending our approach to more general stiff systems or higher-order methods, and further improving gradient approximations in large-scale settings.
@@
Term I is addressed in Corollary S2.9, where we compare the discrete gradient $\nabla_{\Theta}^{\mathsf{p}}E$ (defined in (24)) with the exact gradient $\nabla_{\Theta}E$ (from (15)). These expressions differ only by replacing $\mathsf{p}_{l}$ with $\hat{\mathsf{p}}_{l}$ and $\mathsf{z}_{l}$ with $z(lh)$, and this discrepancy is shown to be $\mathcal{O}(h^{2})$, based on the second-order accuracy and boundedness of the Leapfrog integrator, as established in Lemmas S2.2 and S2.4. The result also relies on the boundedness of $\mathsf{p}_{l}$, $\hat{\mathsf{p}}_{l}$, and the second-order regularity of $f$.
Term II is analyzed in Lemma S2.13, where we compare $L\mathsf{T}\nabla_{\Theta}^{\mathsf{p}}E$ with $L\nabla_{\Theta}^{\mathsf{q}}E$. A direct computation using the structure of the averaging matrix $\mathsf{T}$ (defined in (21)) shows that reordering the averaging introduces only an $\mathcal{O}(h^{2})$ error. The proof uses the boundedness of $\hat{\mathsf{p}}_{l}$, the third-order regularity of $f$, and second-order bounds on $\theta$.
Term III is handled in Corollary S2.11, which compares $\nabla_{\Theta}^{\mathsf{q}}E$ with the functional gradient $\left.\frac{\delta\mathcal{E}}{\delta\theta}\right|_{\theta(t)}(t)$ (defined in (11)). The argument is based on Lemma S2.6, which shows that the averaged quantity $\hat{\mathsf{q}}_{l}$ closely approximates $p(lh)$. Although $\hat{\mathsf{p}}_{l}$ deviates from the Leapfrog update near initialization, averaging corrects this effect. The proof also uses the boundedness of $z$ and the first-order regularity of $f$.
@@
  • A-priori bounds on $p(t)$, $z(t)$, $\mathsf{z}_{l}$, and $\hat{\mathsf{p}}_{l}$, established in Lemma S3.3.
@@
    • Lemma S3.4 and Corollary S3.6 show that $\mathsf{z}_{l}$ and $\mathsf{z}_{l+\alpha}$ approximate $z(lh)$ and $z((l+\alpha)h)$ up to $\mathcal{O}(h^{2})$, due to the second-order accuracy of the 2-stage ERK method.
    • Proposition 2.4 ensures that $\mathsf{p}_{l}$ approximates $p(lh)$ up to $\mathcal{O}(h^{2})$, with supporting bounds from Lemmas S3.9 and S3.11.
    • Corollary S3.13 controls the middle-stage interpolation $\mathsf{q}_{l+\alpha}$, whose accuracy follows from the linear interpolation structure and Proposition 2.4.
@@ Appendix B Experimental details
In this section, we describe more details about our numerical experiments in Section 4. All ODE-nets, including Leapfrog, Midpoint, Ralston, and Nystrom are implemented using PyTorch’s torch.nn module, with gradient directly provided by Pytorch’s autodifferentiation.
@@ B.1 Details of gradient computation experiment
The gradient computation experiments in Section 4.1 considers a simple case, where for all of Leapfrog (Figure 1 and Figure 2), Midpoint (Figure 1 and Figure S1), Ralston (Figure 1 and Figure S2) and Nystrom (Figure 1) we used the initial condition $x=3$, label $y=24$, output function $g$ being the identity function, and the loss function be the mean square loss (14). The initial parameter curve $\theta:[0,1]\to\mathbb{R}^{3}$ is defined to be $\theta(t):=(\frac{t+2}{4},0,1)$, and the activation function is $f(z,\theta):=\theta_{3}\tanh(z\theta_{1}+\theta_{2})$, where $\theta_{1},\theta_{2},\theta_{3}$ denote the first, second and third coordinate of $\theta$ respectively. For simplicity, we only compute and plot the first coordinate of the gradients. We demonstrate the convergence of the $L_{\infty}$ error of our modified gradients using different $L$-s ranging from $4$ to $64$. The gradients in the Euler ODE-Net model are computed using the closed form expression in [9]. In addition, the “ODE-gradients,” which are the values of the functional derivatives $\left.\frac{\delta\mathcal{E}}{\delta\theta}\right|_{\theta(t)}(t)$ at the corresponding time steps, are computed using (11), where for $z(t)$ and $p(t)$ we solve them using Scipy’s scipy.integrate.solve_ivp function.

### v3

@@ 2.3 Oscillations in Auto-Differentiation
To start, we examine Leapfrog. Comparing (11) and (15), we immediately observe some nice features. $\partial_{\theta_{l}}E$ honors the structure of $\frac{\delta\mathcal{E}}{\delta\theta}(lh)$. Both of them see the gradient as a product of $p$, the adjoint vector that runs from final time to $t=0$, and the Jacobian $\partial_{\theta}f$. Moreover, the adjoint vector of $\mathsf{p}$ aligns very well with that of (12). In particular, (16c) is exactly a Leapfrog discretization of (12) for all $l\leq L-3$.
However, one strong drawback of the formula hinges on the definition of $\mathsf{p}_{L-1}$ an $\mathsf{p}_{L-2}$. Intuitively, for a good match, both $\mathsf{p}_{L-1}$ and $\mathsf{p}_{L-2}$ should approximately resemble $p(1)$. Comparing (16a) and (13), however, we see that $\mathsf{p}_{L-1}$ nearly doubles the value of $p(1)$ while $\mathsf{p}_{L-2}$ is $\mathcal{O}(h)$. This immediately introduces $\mathcal{O}(1)$ contrast at the beginning of the evolution of $\mathsf{p}$. This high contrast propagates back in time using the Leapfrog formula (16c), bringing high oscillations throughout. This was seen in Figure 1. Moreover, refining the discretization by enlarging $L$ and reducing $h$ is not helpful. The $\mathcal{O}(1)$ oscillation is encoded in the definition of $\mathsf{p}_{L-1}$ and $\mathsf{p}_{L-2}$, and it does not disappear in the $h\to 0$ limit. The entire derivations are completely exact and honor the computation of auto-differentiation that returns the honest derivatives for the discrete system. These oscillations are not artifacts induced by auto-differentiation itself, but rather the inconsistency embedded in the discrete system.
We now examine ODE-Net generated by ERK-2 stage method. The situation here is even more intricate. This time, one can show that the adjoint vector $\mathsf{p}_{l}$ (18) is actually a second order approximation of $p$ and thus remains faithful to the continuous adjoint $p(t)$.
@@
The proof for this proposition is straightforward, as presented in Supplementary Section SM3.2. Essentially, one needs to define an auxiliary variable, termed $\hat{\mathsf{p}}_{l}$, that solves the same equation (18) with $\mathsf{z}_{l}$ and $\mathsf{z}_{l+\alpha}$ replaced by $z(lh)$ and $z((l+\alpha)h)$ respectively, and deploys the triangular inequality.
@@
Higher order ERK suffers from the same challenge. Though the assembling coefficient vary, the typical situation nevertheless gives incorrect match to its continuous versions, and provides high oscillations. For example, in the case of Nyström (8),
@@ 3 Modified auto-differentiation for ODE-net
The methods provided below require some generic assumptions for the fix to hold. {assumption} The activation function $f(z,\theta)$ satisfies: a) Regularity: $f$ is third order continuously differentiable with derivatives bounded by $C_{a}$; b) Quadratic-linear growth: For all $z\in\mathbb{R}^{d},\theta\in\mathbb{R}^{n}$, $|f|\leq C_{b}(|\theta|^{2}+1)(|z|+1)$ for some $C_{b}$. The output function $g(z)$ to possess: c) Regularity: $g(z)$ and its gradient $\nabla g(z)$ are Lipschitz continuous with Lipschitz constants bounded by $C_{c}$. The training data distribution $\mu$ has: d) Compact Support: There exists some $C_{d}>0$ such that $\text{supp}(\mu)$, the support of $\mu$, lies within a ball centered at the origin with radius $C_{d}$. These assumptions are rather mild and are typically satisfied. For example, the commonly used activation function $f(z;\sigma,W,b)=\sigma\cdot\tanh(Wz+b)$ and the output function $g(z):=z$ satisfy all the conditions above.
@@ 3.1 Modified Leapfrog ODE-net
The core of the proof lies in the realization that “averaging” the gradient is effectively equivalent (up to $\mathcal{O}(h^{2})$ error) to “averaging” the adjoint vector $\mathsf{p}$. The averaging for the gradient is already denoted by $\mathsf{T}$ defined in (21), and we are to define a new matrix $\tilde{\mathsf{T}}$ (see definition in (SM(22))) that acts on $\{\mathsf{p}\}$. Two auxiliary quantities will be defined. Firstly, we define $\hat{\mathsf{p}}=\{\hat{\mathsf{p}}_{l}\}_{l=0}^{L-1}$, a semi-continuous object that uses the same formula (16) with $\mathsf{z}_{l}$ replaced by $z(lh)$. We then define an averaged version:
@@
The theorem is proved if all three terms are $\mathcal{O}(h^{2})$. This can be proved to hold. We discuss the proof strategy in Appendix A.1, and leave details to supplementary materials in Corollary SM2.9, Lemma SM2.13 and Corollary SM2.11 respectively.
@@ 3.2 Modified 2-stage ERK ODE-nets
We should note that $\{\mathsf{p}_{l}\}$ defined by (18) coincide with the output of auto-differentiation $\{\partial_{\mathsf{z}_{l}}E\}$, as justified by the following proposition:
@@
For all $l=0,1,\dots,L$, the solution $\mathsf{p}_{l}$ to  (18) satisfies: $\{\partial_{\mathsf{z}_{l}}E\}$.
The proof is straightforward calculation and is in Supplementary Section SM3.1.
@@
We leave the discussion of the proof strategy to Appendix A.2, with details provided in Supplementary Section SM3. The major component of the proof is Taylor approximation both for the forward and the backward equation to the second order.
@@ 4 Numerical experiments
We present numerical experiments to verify our observations, and validate the correction method we proposed. The implementation details and additional figures are deferred to Appendix B and Supplementary Section SM4 respectively.
Four discretization schemes are studied, and they are Leapfrog (3), an example of LLM, Midpoint and Ralston (6)-(7), as examples of ERK-2 stage, and Nyström (8) as an example of ERK-3 stage respectively. For the first three schemes, we provide corrections.
@@ 4.1 Gradient computation
In Figure 2SM1, and SM2, we plot the performance of auto-differentiation result and the our proposed correction for Leapfrog, Midpoint and Ralston respectively. In each plot, the left panel shows comparison of true gradient (blue), gradient as an output of auto-differentiation (green) and our modified correction (red), following (22) and (26). The right panel demonstrates the convergence rate with respect to $h$. For all three methods, the output of auto-differentiation are highly oscillatory and non-converging: In $L_{\infty}$ norm, the error stays as a constant for all $h$. The modified gradients (red) visually show no difference from the true gradient, and the convergence rate with respect to $h$ are $h^{2.00}$, $h^{1.70}$, and $h^{1.84}$ respectively, confirming the second order accuracy predicted in Theorem 3.1 and 3.4. As a comparison, we also plot out the convergence using the standard forward Euler discretization. The rate stays around $h^{~0.98}$, and thus confirms it as a first order method.
@@ 4.2 Curve-learning
Figure 3: For Leapfrog method. The top row shows the training loss and test error over epochs. The middle row compares $\Theta$ outputs from vanilla (unmodified) auto-differentiation and our modified gradients, and the bottom plot presents the final reconstruction. Quantitatively, the corrected gradients achieve much lower final errors: training loss decreases from $3.0\times 10^{-3}$ to $2.3\times 10^{-4}$, and test error from $3.3\times 10^{-2}$ to $5.4\times 10^{-4}$. These improvements confirm the effectiveness of the proposed correction.
Figure 4: For Leapfrog Method. The figure compares reconstructions of a spiral learned using auto-differentiation gradients (red) and our modified gradients (blue). Our modification achieves results that align more closely with the ground truth (green). Quantitatively, the trajectory Root Mean Square Error (RMSE) drops from $0.75$ to $0.07$, the final-point error from $1.06$ to $0.13$, and the maximum deviation from $1.07$ to $0.15$. These results confirm that the corrected gradients yield trajectories that closely follow the ground-truth dynamics.
@@
We first experimented on learning a linear function. The improvement for Leapfrog, Midpoint and Ralston are respectively shown in Figure 3SM3, and SM4. The top rows show both auto-differentiation updates (in red) and our modified gradients (in blue) decay in error over epochs. In these examples, our test error clearly shows better performance compared to auto-differentiation. The second rows depict the computed $\Theta$. While auto-differentiation leads to strong oscillatory patterns, our modified gradients provide smoother parameters and better convergence in epochs. The last rows present the final function reconstructions. For a linear function, our modified gradients show consistent improvement toward the ground truth (in green).
We also tested learning a spiral in $\mathbb{R}^{2}$. The results for Leapfrog, Midpoint, and Ralston methods are shown in Figures 4, SM5, and SM6, respectively. In each plot, the reconstructed trajectory using auto-differentiation is shown in red, while the reconstruction using the modified gradient is shown in blue. Our modification produces results that more closely match the ground truth, shown in green.
@@ 5 Conclusions
Our findings suggest that careful treatment of discretization and optimization order is crucial for improving stability and accuracy in training physics-inspired neural networks. The proposed post-processing framework in Section 3.2 can in principle be extended to higher-order ERK schemes, including the Nyström scheme illustrated in Figure 1. Under higher regularity assumptions, the auto-differentiated adjoint variables remain high-order approximations of the continuous adjoint, and the main modification lies in constructing higher-order interpolants for intermediate states. While the derivations are algebraically involved, the extension is conceptually straightforward. Future work may explore extending our approach to more general stiff systems or higher-order methods, and further improving gradient approximations in large-scale settings.
@@
Term I is addressed in Corollary SM2.9, where we compare the discrete gradient $\nabla_{\Theta}^{\mathsf{p}}E$ (defined in (24)) with the exact gradient $\nabla_{\Theta}E$ (from (15)). These expressions differ only by replacing $\mathsf{p}_{l}$ with $\hat{\mathsf{p}}_{l}$ and $\mathsf{z}_{l}$ with $z(lh)$, and this discrepancy is shown to be $\mathcal{O}(h^{2})$, based on the second-order accuracy and boundedness of the Leapfrog integrator, as established in Lemmas SM2.2 and SM2.4. The result also relies on the boundedness of $\mathsf{p}_{l}$, $\hat{\mathsf{p}}_{l}$, and the second-order regularity of $f$.
Term II is analyzed in Lemma SM2.13, where we compare $L\mathsf{T}\nabla_{\Theta}^{\mathsf{p}}E$ with $L\nabla_{\Theta}^{\mathsf{q}}E$. A direct computation using the structure of the averaging matrix $\mathsf{T}$ (defined in (21)) shows that reordering the averaging introduces only an $\mathcal{O}(h^{2})$ error. The proof uses the boundedness of $\hat{\mathsf{p}}_{l}$, the third-order regularity of $f$, and second-order bounds on $\theta$.
Term III is handled in Corollary SM2.11, which compares $\nabla_{\Theta}^{\mathsf{q}}E$ with the functional gradient $\left.\frac{\delta\mathcal{E}}{\delta\theta}\right|_{\theta(t)}(t)$ (defined in (11)). The argument is based on Lemma SM2.6, which shows that the averaged quantity $\hat{\mathsf{q}}_{l}$ closely approximates $p(lh)$. Although $\hat{\mathsf{p}}_{l}$ deviates from the Leapfrog update near initialization, averaging corrects this effect. The proof also uses the boundedness of $z$ and the first-order regularity of $f$.
@@
  • A-priori bounds on $p(t)$, $z(t)$, $\mathsf{z}_{l}$, and $\hat{\mathsf{p}}_{l}$, established in Lemma SM3.3.
@@
    • Lemma SM3.4 and Corollary SM3.6 show that $\mathsf{z}_{l}$ and $\mathsf{z}_{l+\alpha}$ approximate $z(lh)$ and $z((l+\alpha)h)$ up to $\mathcal{O}(h^{2})$, due to the second-order accuracy of the 2-stage ERK method.
    • Proposition 2.4 ensures that $\mathsf{p}_{l}$ approximates $p(lh)$ up to $\mathcal{O}(h^{2})$, with supporting bounds from Lemmas SM3.9 and SM3.11.
    • Corollary SM3.13 controls the middle-stage interpolation $\mathsf{q}_{l+\alpha}$, whose accuracy follows from the linear interpolation structure and Proposition 2.4.
@@ Appendix B Experimental details
In this section, we describe more details about our numerical experiments in Section 4. All ODE-nets, including Leapfrog, Midpoint, Ralston, and Nyström are implemented using PyTorch’s torch.nn module, with gradient directly provided by Pytorch’s autodifferentiation.
@@ B.1 Details of gradient computation experiment
The gradient computation experiments in Section 4.1 considers a simple case, where for all of Leapfrog (Figure 1 and Figure 2), Midpoint (Figure 1 and Figure SM1), Ralston (Figure 1 and Figure SM2) and Nyström(Figure 1) we used the initial condition $x=3$, label $y=24$, output function $g$ being the identity function, and the loss function be the mean square loss (14). The initial parameter curve $\theta:[0,1]\to\mathbb{R}^{3}$ is defined to be $\theta(t):=(\frac{t+2}{4},0,1)$, and the activation function is $f(z,\theta):=\theta_{3}\tanh(z\theta_{1}+\theta_{2})$, where $\theta_{1},\theta_{2},\theta_{3}$ denote the first, second and third coordinate of $\theta$ respectively. For simplicity, we only compute and plot the first coordinate of the gradients. We demonstrate the convergence of the $L_{\infty}$ error of our modified gradients using different $L$-s ranging from $4$ to $64$. The gradients in the Euler ODE-Net model are computed using the closed form expression in [DiChLiWr:2022overparameterization, Appendix C.4]. In addition, the “ODE-gradients,” which are the values of the functional derivatives $\left.\frac{\delta\mathcal{E}}{\delta\theta}\right|_{\theta(t)}(t)$ at the corresponding time steps, are computed using (11), where for $z(t)$ and $p(t)$ we solve them using Scipy’s scipy.integrate.solve_ivp function.